# LIR: Efficient Degradation Removal for Lightweight Image Restoration

### 02 Feb 2024 v1

@@ 4.1 Experiment Settings
We set the number of EAA Blocks in the LIR to [3,3,3,3,4], and the width to 48. During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$. Random image cropping and horizontal flipping are used as data augmentation strategies. Following [62], we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising. For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K]. The number of parameters and computation for each method are calculated based on the input size of (3,192,192). Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package. All experiments are conducted using the official release code for each method. For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
Method Type Rain100L [58] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
UMRL [59] CNN 29.18 0.923 0.98 9.31
PReNet [44] RNN 32.44 0.950 0.17 37.27
AirNet [25] CNN 34.90 0.966 8.93 169.47
MPRNet [61] CNN 36.40 0.965 20.13 960.39
SPAIR [42] CNN 36.93 0.969 $\triangle$ $\triangle$
HINet [4] CNN 37.28 0.970 88.67 95.92
MAXIM-2S [53] MLP 38.06 0.977 22.50 80.13
SFNet [7] CNN 38.21 0.974 13.27 70.00
DRCNet [26] CNN 38.23 0.976 18.90 $\triangle$
Restormer [62] Transformer 38.99 0.978 26.13 79.31
LIR (Ours) CNN 38.96 0.983 7.31 63.96
@@ 4.2 Image Deraining
The Rain14000 [15] (11,200 training images) and Rain100H [58] (1,254 training images) are used to train LIR, and both the Rain100L [58] and Rain100H [58] datasets are used for evaluation. The results are presented in Tab. 1 and Tab. 2. In the Rain100L results (Tab. 1), LIR with minimal parameters and computation outperforms the DRCNet [26] and achieves comparable performance to Restormer [62] with only a slight 0.02 dB decrease. As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab. 1. It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours.
In the Rain100H results (Tab. 2), although the number of parameters and computations are reduced by 63% and 93% respectively, the PSNR is still 0.14 dB higher than MPRNet [61]. Furthermore, LIR achieves comparable performance to HINet [4]. Fig. 6 shows visual comparisons among LIR, Restormer, SFNet, and HINet. It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
Method Type Rain100H [58] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
SEMI [56] CNN 16.56 0.486 $\triangle$ $\triangle$
UMRL [60] CNN 26.01 0.832 0.98 9.31
RESCAN [28] CNN 26.36 0.786 0.15 18.07
MSPFN [21] CNN 28.66 0.860 1.65 $\triangle$
MPRNet [61] CNN 30.41 0.890 20.13 960.39
HINet [4] CNN 30.65 0.894 88.67 95.92
LIR (Ours) CNN 30.58 0.904 7.31 63.96
Table 2: Image Deraining results evaluated by PSNR and SSIM on Rain100H dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
Method Type SOTS outdoor [24] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
FD-GAN [10] GAN 23.15 0.920 13.98 38.90
AirNet [25] CNN 23.18 0.900 8.93 169.47
IDR [65] CNN 25.24 0.943 15.34 $\triangle$
GridDehazeNet [32] CNN 30.86 0.982 0.96 12.05
U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$
FFA-Net [43] CNN 33.38 0.980 4.46 161.74
MAXIM-2S [53] MLP 34.19 0.985 22.5 80.13
LIR (Ours) CNN 34.85 0.985 7.31 63.96
@@
Figure 7: The visual comparisons for LIR, Restormer [62], AirNet [25], and MPRNet [61]. LIR produces better visual results.
@@ 4.3 Image Dehazing
The OTS [24] (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor [24] is used for evaluation. In the SOTS outdoor results (Tab. 3), LIR significantly outperforms both the MAXIM-2S [53] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB. Visual comparisons between LIR, FFA-Net, GridDehazeNet, and AirNet are shown in the Fig. 5.
Method Type CBSD68 [37] $\downarrow$ Urban100 [18] Params [M] GFLOPs
$\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
FD-GAN [10] GAN 30.25 28.81 26.43 - - - 13.98 38.90
DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$
MPRNet [61] CNN 33.54 30.89 27.56 - - - 20.13 960.39
AirNet [25] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47
BRDNet [47] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$
IDR [65] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$
DRUNet [67] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78
Restormer [5] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31
LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
@@ 4.4 Image Denoising
We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise. The DFWB (DIV2K [1], Flickr2K, WED [36], BSD500 [2]) datasets with 8,594 training images are used to train LIR. For evaluation, we use the CBSD68 [37] and Urban100 [18] datasets. The light blue color in the Tab. 4 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50). Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training. As shown in the Tab. 4, LIR achieves comparable performance to Restormer [62] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting. Moreover, with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [67] on Urban100. The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig. 7. Although LIR is not as high as Restormer on PSNR, the visual results are far superior.
@@ 4.6 Ablation Studies
We conduct ablation experiments on the Rain100L dataset, using unchanged LIR as the baseline. The ablation experiments are performed on three types: Simple Residual, No Patch Attention, and No Adaptive Filter. Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design. The results in the Tab. 6 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR. In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
Method Type GoPro [38] HIDE [45] Params [M] GFLOPs
PSNR $\downarrow$ SSIM PSNR SSIM
IDR [65] CNN 27.87 0.846 - - 15.34 $\triangle$
DBGAN [66] GAN 31.10 0.942 28.94 0.915 11.58 189.71
MT-RNN [39] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$
DMPHN [63] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$
Zhe et al. [22] CNN 31.79 0.949 - - $\triangle$ $\triangle$
Suin et al. [46] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$
SPAIR [42] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$
LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
Type PSNR Params [M] GFLOPs
Baseline 38.96 7.31 63.96
Simple Residual 38.87 6.86 49.15
No Patch Attention 38.85 6.51 62.15
No Adaptive Filter 38.66 7.31 62.40
@@
5 Conclusion
In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation. Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and a simple structural design. In addition, we propose the Adaptive Filter (AF) to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks. Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab. 6). To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy. Extensive experiments on deraining, dehazing, denoising, and deblurring demonstrate that our LIR achieves comparable performance to state-of-the-art networks with fewer parameters and computations. Notably, our approach achieves state-of-the-art visual results in deraining and denoising tasks.

### 29 Feb 2024 v2

@@ 4.1 Experiment Settings
We set the number of LAA Blocks in the LIR to [3,3,3,3,4], and the width to 48. During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$. Random image cropping and horizontal flipping are used as data augmentation strategies. Following [58], we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising. For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K]. The number of parameters and computation for each method are calculated based on the input size of (3,192,192). Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package. All experiments are conducted using the official release code for each method. For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
[…]
[width=0.6]figures/dehaze.eps
@@ 4.2 Image Deraining
The Rain14000 [15] (11,200 training images) and Rain100H [56] (1,254 training images) are used to train LIR, and the Rain100L [56] datasets is used for evaluation. The results are presented in Tab. 1. In the Rain100L results (Tab. 1), LIR with minimal parameters and computation outperforms the DRCNet [25] and achieves comparable performance to Restormer [58] with only a slight 0.02 dB decrease on PSNR. In addition, our LIR achieves state-of-the-art on SSIM, which significantly outperforms Restormer. As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab. 1.
It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours. Fig. 5 shows visual comparisons among LIR, Restormer, SFNet, and HINet. It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
@@
[…]
0.9 Method Type SOTS outdoor [23] Params [M] GFLOPs PSNR $\downarrow$ SSIM FD-GAN [11] GAN 23.15 0.920 13.98 38.90 AirNet [24] CNN 23.18 0.900 8.93 169.47 IDR [62] CNN 25.24 0.943 15.34 $\triangle$ GridDehazeNet [30] CNN 30.86 0.982 0.96 12.05 U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$ FFA-Net [41] CNN 33.38 0.980 4.46 161.74 MAXIM-2S [52] MLP 34.19 0.985 22.5 80.13 DehazeFormer [44] Transformer 34.95 0.984 2.52 13.37 LIR (Ours) CNN 34.85 0.985 7.31 63.96
[…]
[width=]figures/denoise.eps
Figure 7: The visual comparisons for LIR, Restormer [58], AirNet [24], and MPRNet [59]. LIR produces state-of-the-art visual results.
@@ 4.3 Image Dehazing
The OTS [23] (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor [23] is used for evaluation. In the SOTS outdoor results (Tab. 2), LIR significantly outperforms both the MAXIM-2S [52] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB. In addition, LIR achieves comparable performance on PSNR and better performance on SSIM to DehazeFormer. It is worth noting that GFLOPs and parameters do not accurately reflect the complexity of the transformer model, the actual complexity of LIR is much lower compared to DehazeFormer. Visual comparisons are shown in the Fig. 6.
@@
[…]
0.9 Method Type CBSD68 [35] $\downarrow$ Urban100 [18] Params [M] GFLOPs $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ FD-GAN [11] GAN 30.25 28.81 26.43 - - - 13.98 38.90 DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$ MPRNet [59] CNN 33.54 30.89 27.56 - - - 20.13 960.39 AirNet [24] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47 BRDNet [46] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$ IDR [62] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$ DRUNet [64] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78 Restormer [4] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31 LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
@@ 4.4 Image Denoising
We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise. The DFWB (DIV2K [1], Flickr2K, WED [34], BSD500 [2]) datasets with 8,594 training images are used to train LIR. For evaluation, we use the CBSD68 [35] and Urban100 [18] datasets. The light blue color in the Tab. 3 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50). Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training. As shown in the Tab. 3, LIR achieves comparable performance to Restormer [58] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting. Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [64] on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting. The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig. 7. Although LIR is not as high as Restormer on PSNR, the state-of-the-art visual result is achieved.
[…]
0.9 Method Type GoPro [36] HIDE [43] Params [M] GFLOPs PSNR $\downarrow$ SSIM PSNR $\downarrow$ SSIM IDR [62] CNN 27.87 0.846 - - 15.34 $\triangle$ DBGAN [65] GAN 31.10 0.942 28.94 0.915 11.58 189.71 MT-RNN [37] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$ DMPHN [60] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$ Zhe et al. [21] CNN 31.79 0.949 - - $\triangle$ $\triangle$ Suin et al. [45] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$ SPAIR [40] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$ LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
@@ 4.6 Ablation Studies
We conduct ablation experiments on the Rain100L dataset, using unchanged LIR as the baseline. The ablation experiments are performed on three types: Simple Residual, No Patch Attention, and No Adaptive Filter. Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design. In addition, No Adaptive Filter refers to using a vanilla convolution to replace the Adaptive Filter for fair comparison. Therefore, the number of parameters is unchanged. The results in the Tab. 5 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR. In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
@@
Metric Baseline Simple Residual Patch Attention Adaptive Filter
PSNR 38.96 38.87 38.85 38.66
Params [M] 7.31 6.86 6.51 7.31
GFLOPs 63.96 49.15 62.15 62.40
5 Conclusion and Limitation
5.1 Conclusion
In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation. Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and a simple structural design. In addition, we propose the Adaptive Filter (AF) to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks. Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab. 5). To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy.
5.2 Limitation
For the deraining task, LIR achieves comparable performance with the state-of-the-art method on PSNR and the best performance on SSIM with smaller parameter numbers and computations; For denoising and dehazing tasks, LIR also achieves comparable performance to the state-of-art method. In addition, LIR achieves state-of-the-art visual results on deraining and denoising tasks. However, as Tab. 4 shows, LIR is less prominent on deblurring tasks than deraining and denoising. We believe that this is because, in the deblurring task, object outlines tend to be stretched out very long. While our Adaptive Filter is very effective when dealing with static object contours, it has limitations when dealing with dynamic contours caused by fast-moving objects.